OpenStudy (anonymous):

2^3x = 32 need help please and thanks

OpenStudy (anonymous):

does it help to know that \(32=2^5\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good

OpenStudy (anonymous):

what do i do next

OpenStudy (anonymous):

set \[2^{3x}=2^5\iff 3x=5\]

OpenStudy (pratyush5):



OpenStudy (pratyush5):

3x = 5 is what you get

OpenStudy (anonymous):

x=5/3

OpenStudy (pratyush5):

correct :)